how would i solve for x:
x
___
x - 2 =

4
_____
5

Found 2 solutions by faceoff57, uma:
Answer by faceoff57(108)

x/(x-2) = 4/5
5x = 4(x-2)
5x = 4x - 8
5x - 4x = -8
x = -8

Answer by uma(370)

x/(x-2) = 4/5
Multiplying both the sides of the equation by 5(x-2) we get,
5x = 4(x-2)
==> 5x = 4x - 8
==> 5x - 4x = -8
==> x = -8